A random sample of 9 fields of rye has a mean yield of 37.0 bushels per acre and standard deviation of 6.23 bushels per acre. Determine the 95% confidence interval for the true mean yield. Assume the population is approximately normal.

Step 1 of 2 : Find the critical value that should be used in constructing the confidence interval. Round your answer to three decimal places.
Answer #1

solution: given that, X = 370 S = 6.23 n= g degree of freedom = n-1 = 9-128 At 95% confidence level the tis X=1-95), L=1- 0.9